February 5, 1999 --- SPWG Meeting Minutes

0. Discussion of Gyroless Spacecraft Operations

The first week of March there will be a momentum management, station keeping and roll attitude correction. In the past, star tracker single event upsets (SEUs) caused transition to RMW mode. Now, without gyros: If we are lucky and re-acquire the star quickly after an SEU, we will be in RMW. However, if there is an SEU and the tracker doesn't re-lock quickly, SOHO will go to ESR mode. This issue is being addressed in the gyroless study by MMS.

1. Review of action items

- None Open

2. Boundary conditions

- Upcoming Maneuvers First week in March, is planned for next maneuver operations. - MDI Continuous: MDI Continuous has been moved to avoid maneuvers and roll activities. The three months continuous will now start on March 14. MDI to submit the OCD for this by March 4th. The last month is the best-effort month.

3. Preview of future months

- See Monthly Calendar at

4. Priorities for the coming month

a) Individual instrument plans ------------------------------- TRACE -- Active region studies with VLA on February 13, 15, 19, and 20. Target for 13th will be chosen on 11 February when a representative will come to the EOF. Priorities: SUMER collaboration, flux emergence, flares, high cadence Through 15 Feb: Coordination with Sac Peak observers via Dick Shine. between Feb 8 - 14: Extensive post-eclipse cal.; contact: Ted Tarbell Feb 13, 15, 19, 20: VLA coordination; 4 times 4.0 hrs in DnC configuration and maybe 27 at LST near 0.0; contact (coordinator): Ed DeLuca Feb 16 Lunar occultation Feb 15 - 28 Study of microflares in QS; contact: Arnold Benz and Sam Krucker first-half May: X-ray jets from active regions on the limb; SUMER, TRACE, SXT: contact Davina Innes May 15- June 7: Requested: SVST coordination; contact: Tom Berger June 7-22: SVST coordination; contact Peter Brandt SUMER -- Observing from February 23 - March 23. (Calendar lists planners). Our science plans right now are concentrated around Lyman obervations (Curdt, Heinzel, Tu, Marsch), Ne VIII, Mg X, Fe XII flow and wavelength studies, and spicules for the first week. Depending on the subject, we would also be willing to support JOPs, however, to have control on total count rates etc., request that the SUMER planner be on the JOP team. There will be an important change in SUMER operational constraints, in that we will allow one raster image per study, if this is required as context image. March 4 - 8, Terry Kucera and Art Poland want to concentrate on the flow and flux tubes in and around filaments.. They will get in touch with others regarding collaborations. After March 23, the telemetry is availalable for the taking until sometime in April. The exact date seems to depend on a possible MEDOC campaign. In any case, we will try to observe in May, when Tenerife cooperations are planned, and hopefully in June. CDS -- Plans for CDS include the following: - Week of February 8-14: GIS BLINKER STUDY, CDS and TRACE - CDS/TRACE BLINKER STUDY -- Run quiet Sun blinker study for 8-10 hours. - STREAMER BOUNDARY SCAN - CORONAL HOLE DIAGNOSTICS - CME ONSET CAMPAIGN, Occasional runs as time allows. Run 'old' JOP 67. - WAVE STUDY (A. Fludra) -- at quiet Sun centre run for 8 hour periods MDI support desirable. - TRAPPED WAVES, Yu-Qing Lou (Chicago), A. Poland and A. Fludra Run GMODE (study 198, var.1). Point to the limb in a polar coronal hole. Duration: 24-36 hrs. - SUNSPOT PLUMES, A. Fludra - HIGH CADENCE OBSERVATIONS OF ACTIVE REGION LOOPS, A. Fludra UVCS -- 4 - 15 February -- Instrument tests & CME Watch, L. Gardner 16 - 21 February -- Measurements of Visible-light power spectra L. Ofman 22 - 28 February -- Tests of calibration of white light channel M. Romoli 01 - 28 March -- CME Watch, synoptic, calibration L. Strachan / Dave Rust EIT -- CME watch with 284 (Fe XIV) instead of the usual 195 (Fe XII). One image full-res each 12 minutes while in submode 6. LASCO -- Synoptic observations for next month. Not much till normal roll. C1 testing will continue at a reduced level. Galileo solar conjunction happens 29-Mar. MDI -- Engineering next week; synoptic magnetograms are being generated; also high resolution campaigns to take advantage of new position of the high resolution field-of-view. MDI will have their monthly planning meeting on Tuesday, February 9. b) Intercalibration activities ------------------------------- ICAL001 planned for the week of March 8th - 14th (week 10). c) Joint observing plans ------------------------- - See Monthly Calendar. d) Scheduling SOLs ------------------ Volunteers for several weeks were placed into the calendar.

5. JOPs status

Getting close to 100 JOPs. Paal Brekke has been "volunteered" by those present at the meeting to go through these JOPs and review them. (Paal is not here to defend himself, he'll be in for a surprise when he returns...)

6. New JOPs

- JOP097 -- Karin Muglach OSCILLATIONS IN SUNSPOTS AND ACTIVE REGIONS TRACE, MDI, EIT, Tenerife VTT and GCT, LaPalma DOT and SVST After discussion, this JOP will be placed "on-hold" until the next SPWG. EIT would like to clarify their role, and their is time before it is run in September. - JOP098 -- Neale Ranns Multi-wavelength Observations of Solar Flares. CDS, EIT, MDI, TRACE, Yohkoh This JOP was accepted... will see how it fits in with maneuver dates in early March when maneuver dates are firmly known.

7. AOB

- SWT at ESTEC, February 25 and 26. Let Bernhard know if there are any specific items you want mentioned. There will be an archiving splinter meeting on Friday, February 26. - Next SPWG on Friday, March 26. - Doug Biesecker provided a Whole Sun Month Update. The third installment of the Whole Sun Month Campaigns has been scheduled for June 24th to July 21st. The campaign will be based on the plans and objectives of the second WSM campaign, which had been due to run Aug 1-28, 1998. The second campaign was shortened and reduced because of the loss of signal from SOHO. The third campaign will probably consist of running JOP's 81 (magnetic fields) and 84 (active regions). JOP 77 (coronal velocities) may also be run, though without the participation of LASCO C1. In addition, coronal hole boundary studies and filament channel studies may be conducted. More information on the original plans for the second WSM can be found at As the plans for the third campaign develop and change, they will be posted on the web. - CD-ROMs are being generated by a new system. Starting this summer, the old system will be removed, which means that IF YOU WANT TO REPROCESS ANY DATA FROM BEFORE JANUARY 1999, YOU MUST ASK FOR THE DATA BEFORE JULY 1999. A separate e-mail to the SOHO community will be sent out to the SOC mailing list describing this in detail.